OpenStudy (anonymous):

What does 1 factorial equal?

OpenStudy (zarkon):

1

OpenStudy (anonymous):

To do a factorial, multiply all the numbers between that number and 1.

OpenStudy (anonymous):

4! = 4*3*2*1 3! = 3*2*1 2! = 2*1 1! = 1

OpenStudy (anonymous):

Ok thanks... I was just confused bc I thought it could be zero

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

The stranger thing is that 0! is also 1.

OpenStudy (anonymous):

yeah